OpenStudy (anonymous):

The lengths of two sides of a right triangle are given. Find the length of the third side. Round to the nearest tenth if necessary. lengths: 28 in. and 15 in. 37.5 in. 23.6 in. 29.6 in. 31.8 in.

OpenStudy (anonymous):

@ryanwhite499

OpenStudy (ryanwhite499):

What do you think?

OpenStudy (anonymous):

I'm doing a pretest so I don't know

OpenStudy (ryanwhite499):

=sqrt(28^2+15^2)

OpenStudy (ryanwhite499):

x^2 = 28^2 + 15^2 = 1009 x= sqrt(1009) = 31.8 in

OpenStudy (ryanwhite499):

Make sure you medal me